OpenStudy (anonymous):

the 20th root of(-8)^20

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (angela210793):

=-8

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

its \[-8\]

OpenStudy (anonymous):

say -8 is x so (-8)^20 = x^20 therefore the 20th root of x^20 = x which is -8

OpenStudy (anonymous):

I am not too sure but i somehow think it is +8 and not -8 bcoz say you take square root of(-8)^2 the root will be +8 and not -8 it should be applicable here also so i think it is +8 not -8. (sqrt(x^2)=|x|)

OpenStudy (angela210793):

yea ur right Shankvee but isn't it \[(\sqrt[20]{-8})^{20}\]

OpenStudy (anonymous):

20 root of -8 results in a complex number.

OpenStudy (anonymous):

thanks so much